Circular No. 5761

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 A. V. Filippenko and T. Matheson, University of California, Berkeley,

communicate: “The brightness of GRO J0422+32 appears to have dropped precipitously during the past month, after having declined slowly prior to this. Very rough estimates made on Apr. 14 UT with the Shane 3-m reflector at Lick Observatory yield V = 17.5-18, as compared with V = 14.6 on Mar. 13 (IAUC 5734). The spectrum (range 300-1000 nm), although noisy, seems rather featureless except for prominent Hα emission.”

SUPERNOVA 1993J IN NGC 3031 M. Centurion, Instituto de Astrofisica de Canarias; and G. Vladilo,

Osservatorio Astronomico di Trieste, write: “During Apr. 6-9, while observing SN 1993J at high spectral resolution (FWHM = 0.015 nm) with the Utrecht échelle spectrograph on the William Herschel telescope at La Palma, we detected the K I (neutral potassium) doublet at 767.6 nm, finding K I absorptions at heliocentric velocities of –6 (EW = 0.004 nm), +118 (0.001 nm), and +135 km/s (0.001 nm), in agreement with the velocities of three of the strongest Na I components reported by King (IAUC 5755).”

D. R. Zurek, R. M. Robb, and D. D. Balam, University of Victoria, report: “We have obtained BVRI observations of SN 1993J using the 0.5-m reflector (+ CCD) as follows: Apr. 12.22 UT, V = 11.18 ± 0.05, B–V = +0.53, V–R = +0.35, V–I = +0.48; 13.18, 11.01 ± 0.05, +0.49, +0.35, +0.35. Re-analysis of photometric data reported on IAUC 5731 using the photometric standards of Corwin (IAUC 5742) yields V = 10.74 ± 0.05, B–V = –0.11, V–R = +0.09, V–I = +0.10 on Mar. 30.25. Photometry in R and I obtained on Apr. 12.22 shows that the supernova has reached a brightness consistent with that obtained on Mar. 30.25.”

SUPERNOVA 1955S IN UGC 9933 Jean Mueller reports her discovery of an apparent supernova on Palo-

mar Sky Survey plates taken 1955 Apr. 1 by G. Abell. SN 1955S was then at mag ∼ 17.5 and located 17′′ west and 17′′ north of the center of UGC 9933 (α = 15^{h}35^{m}.0, δ = +43^{o}43′, equinox 1950.0). There is no object at this position on a J plate taken 1993 Mar. 21 by C. Brewer and J. D. Mendenhall in the course of the second Palomar Sky Survey.

1993 April 14 Daniel W. E. Green